A pool measuring 10 meters by 20 meters is surrounded by a path of uniform width. If the area of the pool and the path combined is 600 square meters, what is the width of the path?
Thank you

Answer by khwang(438)

Let the width of the path be x meters,the combined width and length would
be 10+2x and 20+2x.
So we have
(10 + 2x) (20+2x) = 600,
Divided both sides by 4:
(x+5)(x + 10) = 150,
Multiply out the left hand side and simplify:
x^2+ 15 x -100 = 0,
Factor:
(x+20) (x-5) = 0
So, x = 5 or x = -20 (negative,invalid)
Answer: 5 meters.